OpenStudy (aaronandyson):

calculate the area and the peremeter of the remaining portion,after cutting four quadrants , each radius 7cm,from the four corners of a rectangle measuring 20 cm by 30 cm

OpenStudy (aaronandyson):

@butterflydreamer @ganeshie8 @perl @Sadworld

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (rational):

draw it out

OpenStudy (aaronandyson):

idk...

OpenStudy (rational):

this is ur problem so you should know what the problem is about atleast

OpenStudy (aaronandyson):

idk how to draw thats it

OpenStudy (rational):

read the problem 2-3 times

OpenStudy (aaronandyson):

i did cant understand

OpenStudy (aaronandyson):

shall i calculate the area of the rectangle?

OpenStudy (rational):

start by drawing this first ` rectangle measuring 20 cm by 30 cm`

OpenStudy (rational):

always draw a rough diagram first for geometry problems even if it is not necessary

OpenStudy (rational):

that helps in understanding the problem better

OpenStudy (aaronandyson):

20 is l and 30 is w?

OpenStudy (rational):

looks like it

OpenStudy (aaronandyson):

i drew the rectangle on my paper now?

OpenStudy (rational):

after cutting 7cm radius pieces from the corners, it should look like below : http://community.coreldraw.com/cfs-file/__key/telligent-evolution-components-attachments/00-246-01-00-00-03-42-07/scalloped-rectangle.jpg

OpenStudy (aaronandyson):

okay , now?

OpenStudy (rational):

label the lengths and see if you can find the area and perimeter

OpenStudy (rational):

let me know if you get stuck at a specific step

OpenStudy (aaronandyson):

p is 100cm? a is 600cm^2?

OpenStudy (rational):

those are for rectangle with out cutting the corners

OpenStudy (rational):

perimeter and area will change if you remove the corners

OpenStudy (aaronandyson):

???????how do i calculate then????????????

OpenStudy (rational):

for area, you may subtract the removed area from the total area

OpenStudy (rational):

whats the area of circle of radius 7cm ?

OpenStudy (aaronandyson):

quarter circle or a full circle which one?

OpenStudy (rational):

full circle

OpenStudy (rational):

four quarter circles make a full circle

OpenStudy (aaronandyson):

after cutting the circle is it 446cm^2?

OpenStudy (aaronandyson):

@rational ?

OpenStudy (rational):

are you using 3.14 for pi ?

OpenStudy (rational):

show ur work

OpenStudy (aaronandyson):

before cutting rectangle = 20 times 30 = 600 circle = pi time r^2 = 22/7 times 7 times 7 = 154 sub them = 600-154 = 446

OpenStudy (aaronandyson):

sub them = 600-154 = 446 after cutting the circle ^^^

OpenStudy (rational):

that looks perfect for area ! il let you work the perimeter

OpenStudy (aaronandyson):

p = 2l + 2w = 40 + 60 = 100cm

OpenStudy (aaronandyson):

for the rectangle

OpenStudy (aaronandyson):

p = 2*pi*r = 2 times 22/7 time 7 = 44cm for the circle

OpenStudy (aaronandyson):

sub them ans you get 56

OpenStudy (aaronandyson):

i'm right?

OpenStudy (rational):

you need to be more careful with perimeter notice that you need to add the perimeter of circle here because you're creating extra curves

OpenStudy (aaronandyson):

?????????????????????????????????????????????????????????

OpenStudy (rational):

add the perimeter of circle and subtract the removed edges

OpenStudy (rational):

you're removing two edges of 7cm length from each corner so in total you're removing 2*7*4 = 56cm from the perimeter

OpenStudy (rational):

you're replacing that 54cm perimeter with a circle whose perimeter is 2*pi*7 = 44cm

OpenStudy (rational):

so the perimeter of this rectangle with curved corners would be : \[100 +44 - 56\]

OpenStudy (rational):

simplify